A bullet of mass 20 g, travelling at a speed of 350 m/s, strikes a steel plates at an angle of 30 degree with the plane of plate. it ricochets off at the same angle at a speed of 320 m/s. What is the magnitude of the impulse that the steel plate gives to the projectile ?If the collision with the plate take place over a time delta t=1/1000 s, what is the average force exerted by the plate on the bullet?
1
Expert's answer
2017-03-08T13:34:06-0500
The impulse conservation gives m*v1*sina = p - m*v2*sina. Thus, p = m*sina*(v1 + v2) = 6.7 kg*m/s or F = p/t = 6.7/0.001 = 6700 (N)